### v2

# Disentangling Core and Edge Mechanisms of the L-mode Density Limit in DIII-D

Abstract
The L-mode density limit is re-examined on the DIII-D tokamak using a negative triangularity configuration to study the underlying physics. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit ($n_{\mathrm{G}}$) until the termination of auxiliary heating. Systematic power scans reveal distinct power scalings for the core ($n_{e}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) and edge ($n_{e}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$) density limits. The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below $n_{\mathrm{G}}$. In contrast, the core density continues to rise until it saturates at a level featured by substantially enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis $\approx 6$), elevated Hurst exponents, and a $1/f$-type power spectrum. These findings demonstrate that the L-mode density limit is determined by a combination of edge radiative instabilities and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
I Introduction
High plasma density is a fundamental requirement for achieving a self-sustaining fusion reaction, as fusion power scales strongly with density [1]. However, tokamak operation is constrained by an operational density limit [2, 3, 4]. Exceeding this critical density leads to a rapid loss of confinement, often culminating a plasma termination. Overcoming this long-standing obstacle is therefore a critical step toward realizing economically viable fusion energy.
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: $n_{\text{G}}\,(10^{20}\,\mathrm{m^{-3}})=\frac{I_{\text{p}}\,(\text{MA})}{\pi a^{2}\,(\mathrm{m^{2}})}$, where $I_{\text{p}}$ denotes the plasma current and $a$ the minor radius. This scaling is notably independent of heating power. This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power, particularly in the low-confinement (L-mode) regime. Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that even in L-mode, the achievable density limit may scale positively with increased heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physics based on power balance that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
The underlying physics of the L-mode density limit involves a complex interplay of several processes. Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled synergistically [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the L-H transition, edge transport barriers (ETBs), and edge localized modes (ELMs), which are characteristic of conventional H-mode plasmas. The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of a strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
This paper presents a re-examination of the underlying physics governing the L-mode density limit using NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core ($n_{e}\propto P^{0.27\pm 0.03}_{\text{SOL}}$) and the separatrix ($n_{e}\propto P^{0.42\pm 0.04}_{\text{SOL}}$) regions. Our detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained substantially by enhanced turbulence driven by radiative instabilities and edge cooling. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of a single disruptive limit, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
The remainder of this paper is structured as follows. Section II describes the experimental arrangement, including the plasma configuration and the diagnostic suite employed in this study. Section III presents our experimental findings, organized into four subsections: (1) high-density operation beyond the Greenwald density and the power dependence of the core and edge density limits; (2) the temporal evolution of equilibrium profiles and turbulence approaching the density limit; (3) the edge density limit and the role of radiative edge cooling; (4) the core density limit and the dominant role of avalanche-like turbulence. Finally, Section IV summarizes the key findings and discusses their broader implications for fusion research.
II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power varied provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion $\mathbf{B}\times\nabla\mathbf{B}$ drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ($\delta_{\text{top}}=-0.3$ and $\delta_{\text{bot}}=-0.62$), moderate elongation ($\kappa\approx 1.3$), and an edge safety factor of $q_{95}=4.1$. This paper focuses specifically on discharges with toroidal field $B_{t}=2.0$ T and plasma current $I_{p}=0.6$ MA. Additional experiments exploring different plasma current and toroidal field combinations will be reported separately [26]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
Figure 1: Cross-section of DIII-D tokamak showing locations of key diagnostics used for the present study.
A comprehensive suite of diagnostics systems monitored the evolving plasma state, including its equilibrium profiles, radiated power, and turbulent fluctuations. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [27]. Charge exchange recombination (CER) spectroscopy [28] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles ($E_{r}$) using the carbon ion force balance equation. Bolometer arrays [29] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along a horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side). The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of high-sensitivity beam emission spectroscopy (BES) [32] measured ion-scale ($k_{\theta}<3\,\mathrm{cm^{-1}}$) density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [33, 34] measured turbulent flows and low-to-intermediate-$k$ ($2<k_{\theta}<10\,\mathrm{cm^{-1}}$) density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [35].
III Results
III.1 Achieving High Density Beyond the Greenwald Limit
III.1.1 Density Saturation Beyond the Greenwald Limit
An important result of this study is demonstration of the sustained L-mode plasma operation at densities exceeding the conventional Greenwald density limit. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at $\rho\approx 0.7$, which is similar to the line-averaged density, consistently exceeds the Greenwald density ($n_{\text{G}}$). Crucially, the discharges with medium to high power NBI ($P_{\text{NBI}}>4$ MW) did not disrupt upon reaching their maximum density. Instead, they entered a phase of sustained saturated density that persisted until the termination of auxiliary heating.
Figure 2: Time history of relevant plasma parameters in two representative NT discharges with different power input levels. (a) NBI power input, (b) core density ($\rho=0.7$), (c) separatrix density, (d) the radiated power in the divertor, and (e) radiated power in the core or mainchamber. The dashed horizontal line in (b) and (c) indicates the Greenwald density.
The radiated power in the divertor (Fig. 2(d)) and the main chamber (Fig. 2(e)) also increase with density initially, indicating enhanced radiative losses. The spikes in the divertor radiation power (Fig. 2(d)) are associated with the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge. For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26]. This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density that is presented during high-density operation. Therefore, Thomson scattering measurements are used for the analysis in this study.
III.1.2 Distinct Power Scaling for Core and Edge Density Limit
To explore the relationship between input power and achievable density, we examine the dependence of the maximum density on the power flowing into the scrape-off layer (SOL), $P_{\mathrm{SOL}}$. Here, $P_{\mathrm{SOL}}=P_{\text{in}}-P_{\text{rad,core}}-\frac{\mathrm{d}W}{\mathrm{dt}}$, accounting for total input power, core radiated power, and changes in stored energy. Averaged values over a 100 ms window are used.
Figure 3: The maximum density achieved at two different radial positions (core and separatrix) in negative triangularity (NT) discharges as a function of the loss power entering the scrape-off layer ($P_{\mathrm{SOL}}$). The horizontal dashed line indicates the Greenwald density.
Figure 3 shows different scaling relationships for the core ($\rho=0.7$) and at the separatrix. The density at $\rho=0.7$ is chosen to represent the core region, as it is close to the line-averaged density measured by the Thomson scattering. The maximum core density increases with a relatively weak power dependence, scaling as $n_{e,\text{core}}\propto P^{0.27\pm 0.03}_{\mathrm{SOL}}$. This is consistent with L-mode density limit scalings observed on other devices such as ASDEX Upgrade [8]. In contrast, the density near the separatrix exhibits a stronger dependence, scaling as $n_{e,\text{sep}}\propto P^{0.42\pm 0.04}_{\mathrm{SOL}}$. This power dependence is reminiscent of the power-dependent density limit observed in stellarators [9, 10], suggesting a more universal physical basis involving power balance and transport.
These disparate scalings are a key finding, as they imply that distinct physical processes govern density saturation in the core and edge. The remainder of this paper will investigate these processes, proposing that turbulence dominates the core density limit, while radiative dynamics are critical at the edge.
III.2 Temporal Evolution of Profiles Approaching the Density Limit
III.2.1 Profiles Evolution Towards Density Limit
Figure 4: Temporal evolution of equilibrium profiles in a high-power discharge (shot 194133). (a) Electron density $n_{e}$, (b) electron temperature $T_{e}$, (c) carbon ion temperature $T_{i}$, and (d) toroidal rotation rate $\omega_{\phi}$. The profiles are colored according to the time in milliseconds (ms), with earlier times shown in yellow and later times shown in purple. The dashed vertical line indicates the location of the separatrix.
To investigate the underlying physical mechanisms, the temporal evolution of key plasma profiles is examined as each discharge approaches the density limit. Figure 4 displays the evolution of electron density ($n_{e}$), electron and ion temperature ($T_{e}$, $T_{i}$), and toroidal rotation rate ($\omega_{\phi}$) for a high-power discharge. As the overall density is ramped up (indicated by the color change from yellow to purple), several critical changes emerge.
While the density profile (Fig. 4(a)) rises across the plasma radius, it remains relatively flat in the core, indicating persistent particle transport. Concurrently, both the electron and ion temperature profiles (Figs. 4(b) and 4(c)) decrease substantially. A prominent feature is the flattening of the $T_{e}$ profile at the edge, a classic sign of intensified edge cooling. The toroidal rotation profile (Fig. 4(d)), which is initially peaked in the core, collapses as the density increases, leading to a significant reduction in core velocity shear.
Figure 5: The temporal evolution of the edge radial electric field ($E_{r}$) at the outer mid-plane in a high-power discharge (shot 194133). The profiles are color-coded by time, with earlier times shown in yellow and later times in purple. The vertical dashed lines indicate the approximate location of the separatrix.
Figure 5 shows the evolution of the edge radial electric field profile ($E_{r}$) at the outer mid-plane in the high-power discharge, derived from the carbon ion force balance equation. The $E_{r}$ profile initially peaks in the core region ($\rho<0.8$), but collapses progressively as the density increases. The edge $E_{r}$ profile remains relatively flat throughout the density ramp-up. This weakening of $E_{r}$ is primarily attributed to the reduced core toroidal rotation discussed previously.
These profile changes point towards a degradation in plasma confinement. The collapse of the core rotation and the associated $E_{r}\times B$ velocity shear likely weakens the turbulence suppression. This loss of flow shear facilitates the enhanced turbulent transport that ultimately limits the core density, a topic explored in subsequent sections.
III.2.2 Divergent Evolution of Turbulence in Different Regions
Figure 6: Time evolution of electron density and temperature and turbulence characteristics during a high-density discharge (shot 194133): (a) Local $n_{e}$ measured at $\rho$ = 0.7 and 0.95; (b) Local $T_{e}$ at $\rho$ = 0.7 and 0.95; (c) Spectrogram and (d) RMS amplitude of density fluctuations measured by Doppler back-scattering at the low-field side mid-plane ($\rho\approx 0.95$); (e) Spectrogram and (f) RMS amplitude of line-integrated density fluctuations measured by CO2 interferometer at the high field side; (g) Spectrogram and (h) RMS amplitude of line-integrated density fluctuations measured by RIP along the mid-plane. The vertical dashed line indicates the onset of enhanced turbulence.
As the plasma approaches density saturation, a clear divergence in turbulent behavior emerges between different plasma regions. Figure 6 illustrates this by tracking electron density and temperature alongside multichannel turbulence measurements during a high-power discharge. The vertical dashed line at approximately 3700 ms marks the onset of this divergence, which features enhanced turbulence at the HFS and the core, likely due to radiative cooling. This critical event triggers distinct turbulence responses.
Initially, the local density increases across the radius (Fig. 6(a)). Concurrently, the local intermediate-$k$ turbulence at the low-field side (LFS) mid-plane ($\rho\approx 0.95$), measured by the Doppler backscattering systems, shows a steady increase in the root-mean-square (RMS) amplitude (Figs. 6(e) and 6(f)). However, following the onset of the radiative cooling, a dramatic shift occurs: the edge density abruptly saturates (blue in Fig. 6(a)), the electron temperature drops sharply (Fig. 6(b)), and the LFS edge turbulence amplitude also saturates (Fig. 6(d)).
In contrast, line-integrated, low-$k$ density fluctuations at the HFS (Figs. 6(e) and 6(f)) and midplane (Figs. 6(g) and 6(h)) show a different evolution. The fluctuation amplitudes from both the HFS CO2 interferometer and the midplane RIP diagnostic begin to grow substantially after the onset of edge cooling. This divergent behavior—saturation of localized LFS turbulence contrasted with a continued, strong rise in line-integrated HFS and midplane fluctuations—points to a significant poloidal asymmetry in the transport. The turbulence becomes intense in the HFS region where the radiative front develops, linking the enhanced fluctuations and transport to the dynamics of the edge cooling potentially driven by radiative instabilities.
III.3 The Edge Density Limit: Role of Radiative Cooling and Instabilities
III.3.1 Development of Radiation Front at the HFS edge
The evolution of profiles and turbulence suggests the formation of a radiative front and the resulting edge cooling as a potential limiting mechanism for the edge density. To elucidate the role of this radiative instability, the analysis now focuses on the edge region to examine the dynamics of the 2D radiated power distribution.
Figure 7: Temporal evolution of the 2D distribution of radiated power during a representative high-power discharge (shot 194133), reconstructed from bolometer measurements. (a) Before density ramp-up ($t=2800$ ms), (b) enhanced divertor radiation ($t=3800$ ms), (c) radiation front leaving inner divertor ($t=4800$ ms), and (d) radiation front intensified at the HFS ($t=5800$ ms).
Figure 7 shows the temporal evolution of the radiated power distribution, reconstructed from bolometer measurements in a high-power discharge. This radiative feature is geometrically similar to a MARFE (Multifaceted Asymmetric Radiation From the Edge) [16, 36]. However, in these experiments, it manifests as a non-disruptive, quasi-stable radiation front that leads to enhanced fluctuations and density saturation. The intense but localized radiative edge cooling has also been observed in divertor detachment experiments in NT discharges [37].
Initially (Fig. 7(a), $t=2800$ ms), before the density ramp-up, radiation is concentrated near the inner divertor target, typical of an attached divertor regime. As the density is raised (Fig. 7(b), $t=3800$ ms), enhanced radiation emerges near the X-point, signaling the development of the radiative front. With a further density increase (Fig. 7(c), $t=4800$ ms), the condensation front becomes fully developed and moves upward along the high-field side separatrix, a process strongly correlated with the edge cooling shown previously in Fig. 6. Finally, the radiation front propagates further towards the HFS (inboard), and the radiated power near the separatrix intensifies (Fig. 7(d), $t=5800$ ms). It is noteworthy that in this discharge the disruption occurs shortly after a drop in the NBI power (at $t\approx 5800$ ms), which triggered an $n=1$ locked mode, and is not directly caused by the radiative edge cooling.
III.3.2 Edge Density Saturation by Radiative Instabilities
Previous observations suggest that the edge density saturates as radiative cooling intensifies. To quantify this relationship, we compare the evolution of the edge density with the amplitude of line-integrated density fluctuations. As the interferometer chord on the high-field side (HFS) passes directly through the developing radiation front, its signal serves as a reliable proxy for the intensity of the radiative instability.
Figure 8: Edge density at $\rho=0.95$ is plotted against the RMS amplitude of the density fluctuations measured by the HFS CO2 interferometer. The power law relationships for low- and high-turbulence regimes are indicated by the gray and cyan lines, respectively. The plot is in log-log scale. The data points are colored according to the energy confinement factor $H_{98y2}$. The dashed horizontal line represents the Greenwald density.
Figure 8 plots the edge density (at $\rho=0.95$) against the RMS amplitude of density fluctuations on a log-log scale. The data are collected from multiple discharges and colored by the energy confinement factor $H_{98y2}$, reveal a sharp bifurcation between two distinct operational regimes.
In the low-turbulence regime, characterized by high $H_{98y2}$ values, the edge density increases nearly linearly with HFS turbulence, following a power law of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{1.01}$ (gray line in Fig. 8(a)). This indicates a phase where the increasing density turbulence does not yet limit the achievable edge density and plasma confinement.
This relationship breaks down abruptly as the HFS turbulence exceeds a critical threshold, marking a transition to a high-turbulence regime. Here, the fluctuation amplitude, now dominated by the radiative instability itself, continues to grow by nearly an order of magnitude. In stark contrast, the edge density saturates completely, exhibiting a nearly flat scaling of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.1}$ (cyan line in Fig. 8(b)). Critically, this density saturation is accompanied by a sharp degradation in confinement, with $H_{98y2}$ dropping significantly. The overall scaling relationship is $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.26}$. This analysis suggests that the edge density limit is not a gradual process, but a sharp transition triggered by enhanced turbulence, likely driven by a radiative instability and the associated edge cooling. This instability clamps the edge density at a level consistently below the Greenwald limit and can be related to the loss of the edge confinement.
III.3.3 Turbulence Dynamics and the Breakdown of Zonal Flow Regulation
While the preceding analysis points to the radiative instability as the primary limit on edge density, it is crucial to understand the interplay between the turbulence and the developing radiation front. For this purpose, we use 2D measurements from the BES diagnostic, which provides a detailed look at ion-scale density fluctuations. This analysis utilizes a lower-power discharge, where the radiative cooling occurs at lower density, allowing for higher-quality BES signals.
Figure 9: Radial profiles of LFS edge turbulence characteristics during radiative cooling in a low-power NT discharge (shot 194131): (a) adiabaticity parameter, (b) normalized density fluctuation amplitude, (c) mean poloidal phase velocity, and (d) Reynolds stress $\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, (e) particle flux $\langle\tilde{n}\tilde{v}_{r}\rangle$, and (f) turbulent spreading flux $\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$. The profiles are colored by the time from lighter to darker. Particularly, the profiles corresponding to the onset of the radiative cooling (about 3700 ms) are marked by dashed curves in purple.
Figure 9 shows the evolution of key turbulence characteristics at the LFS edge as the radiative front forms (onset around 3700 ms, dashed purple curves). A fundamental change in the turbulence character is signaled by the dimensionless adiabaticity parameter [38, 13], $\alpha_{\text{adia}}=k_{\parallel}^{2}v_{te}^{2}/(\omega\nu_{ei})$. Here, $k_{\parallel}=1/qR$ is the parallel wavenumber, $v_{te}$ is the electron thermal velocity, $\nu_{ei}=\frac{\sqrt{2}e^{4}n_{i}\ln\Lambda}{12\pi^{3/2}\epsilon_{0}^{2}m_{e}^{1/2}T_{e}^{3/2}}$ is the electron-ion collision frequency, and $\omega$ is the turbulence frequency. As the edge cools, $\alpha_{\text{adia}}$ drops below unity, indicating a transition to a non-adiabatic regime, which is known to enhance turbulence and weaken zonal flow generation [13, 14].
Consistent with this transition, the normalized density fluctuation amplitude ($\tilde{n}/\bar{n}$, Fig. 9(b)) grows steadily. The mean poloidal phase velocity ($\bar{V}_{\theta}$, Fig. 9(c)) remains relatively unchanged until the onset of radiative cooling. Crucially, the ability of the turbulence to self-regulate via zonal flows collapses. This is evident in the Reynolds stress ($\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig. 9(d)), the primary drive for zonal flows. The Reynolds stress flattens, indicating that the nonlinear transfer of energy from micro-turbulence to stabilizing macroscopic flows, $\bar{V}_{\theta}\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, has been severely compromised.
The consequence of this weakly regulated turbulence is an increase in outward transport. The turbulent particle flux ($\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig. 9(e)) is directed outward and grows, while the turbulent spreading flux ($\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig. 9(f)) indicates that turbulence intensity is spreading inward from the separatrix. These detailed measurements indicate a potential causal sequence: edge cooling drives a transition to a non-adiabatic turbulent state where the zonal flow self-regulation mechanism fails. This leads to enhanced outward transport, providing the micro-physical link between the onset of strong radiative cooling and the ultimate saturation of the edge density.
III.4 The Core Density Limit: Role of Avalanche-like Turbulence
While the edge radiative condensation directly limits the edge density, the core density continues to increase, even surpassing the Greenwald density after the radiative front has formed. This divergence strongly suggests that a different, core-localized mechanism is at play. We now present evidence that this mechanism is a fundamental change in the nature of core turbulent transport.
III.4.1 Signatures of Avalanche-like Transport in the Core
The collapse of stabilizing $E_{r}\times B$ shear (as shown in Fig. 5) allows enhanced core turbulence. This is confirmed by examining line-integrated density fluctuations from the RIP diagnostic at the midplane. As the plasma approaches its density limit, the statistical character of this turbulence undergoes a significant change.
Figure 10: (a) Time evolution of the probability density function of the amplitudes of line-integrated density fluctuations measured by the RIP system at the midplane. (b) The time history of the kurtosis of the amplitudes of the line-integrated density fluctuations. The horizontal dashed line presents the kurtosis value of a Gaussian distribution.
Figure 10(a) shows the evolution of the probability density function (PDF) of these fluctuation amplitudes (high-pass filtered at $f>5$ kHz to avoid machine vibration noise). Early in the density ramp-up, the PDF is Gaussian (yellow), characteristic of diffusive, random-walk-like transport. However, as the density limit is approached—and particularly after the onset of the edge radiative cooling—the PDF develops pronounced “heavy tails” (purple). This indicates a transition to a state dominated by large-amplitude, intermittent transport events, a clear departure from simple diffusive behavior.
This transition is quantified by the kurtosis of the PDF (Fig. 10(b)), a measure of its “tailedness.” The kurtosis rises significantly after the onset of the radiative cooling and continues to increase, confirming the growing intermittency. Further evidence comes from the power spectrum of the fluctuations (Fig. 11(a)). After the radiative cooling onset, a distinctive “$1/f$-type” power-law scaling emerges at low frequencies. This type of spectrum is a recognized signature of self-organized criticality (SOC) and avalanche-like dynamics [39, 40, 41], where small perturbations can trigger cascading transport events of all sizes.
Figure 11: (a) Auto-power spectrum of the line-integrated midplane density fluctuations measured by RIP at different times. The dashed line indicates the power-law spectrum of $1/f^{0.8}$. (b) Time evolution of the Hurst exponent calculated from the RIP data (blue) and energy confinement factor (orange). The horizontal dashed line indicates a Hurst exponent of 0.5, corresponding to a random walk process.
Finally, the transition to an avalanche-like state is supported by the Hurst exponent ($H$), a measure of long-range temporal correlations in the data (Fig. 11(b)). The Hurst exponent rises from $H\approx 0.5$ (indicative of uncorrelated, random events) to $H\approx 0.8$ (indicative of persistent, correlated dynamics) as the density limit is approached. Consistently, the energy confinement factor $H_{98y2}$ also decreases substantially when the Hurst exponent $H$ begins to increase, indicating a loss of confinement associated with this transition. Collectively, the heavy-tailed PDFs, increased kurtosis, $1/f$ spectrum, and elevated Hurst exponent point to a consistent picture: core transport transitions from a diffusive to an intermittent, avalanche-like regime, which is ultimately responsible for limiting the core density.
III.4.2 Core Density Saturation by Avalanche-like Turbulence
The evidence for a transition to avalanche-like transport leads to a clear hypothesis: this enhanced turbulence is the primary factor limiting the core density. To test this hypothesis, we examine the direct relationship between the core density (at $\rho\approx 0.7$) and the amplitude of the midplane turbulence, as shown in Fig. 12. The data from multiple discharges, plotted on a log-log scale and colored by the energy confinement factor ($H_{98y2}$), reveal a distinct bifurcation between two transport regimes.
In the low-turbulence regime, corresponding to high confinement ($H_{98y2}>1.0$, yellow data points), the core density rises robustly with the turbulence amplitude, following a relatively strong power-law scaling of $n_{\text{e,core}}\propto\tilde{n}_{rms}^{0.49}$. This represents a phase where the core can still be effectively fueled, and the density increases despite the presence of turbulence.
Figure 12: Core density at $\rho=0.7$ is plotted against the RMS amplitude of the density fluctuations measured by the midplane RIP interferometers on a log-log scale (a) and linear scale (b). The data points are colored according to the normalized total radiation. The dotted horizontal line represents the Greenwald density. Power law relationships for low- and high-turbulence regimes are indicated by the gray and cyan lines, respectively.
However, as the avalanche-like transport intensifies and confinement degrades ($H_{98y2}<1.0$, purple data points), this relationship changes dramatically. The scaling becomes substantially weaker, flattening to $n_{\text{e,core}}\propto\tilde{n}_{\text{rms}}^{0.17}$. This transition to a weaker scaling indicates that the avalanche-like density fluctuations increase turbulent transport so effectively that they prevent significant increases in the core density. Notably, this saturation occurs at densities well above the Greenwald limit, reaching nearly twice the Greenwald density.
The saturation of the core density is more clearly visible in the linear-scale plot in Fig. 12(b). The overall response of the core density to the turbulence amplitude scales as $n_{\text{core}}\propto\tilde{n}_{\text{rms}}^{0.26}$, which also underscores the saturated behavior with increasing turbulence. The weak dependence on turbulence in this regime aligns well with the macroscopic power scaling of the core density limit ($n_{\mathrm{e,core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found earlier. This alignment, combined with the avalanche-like dynamics, suggests that the core density limit is governed by a fundamental transition in the core turbulent transport itself, which ultimately saturates the density profile.
IV Conclusion and Discussion
This study has re-examined the L-mode density limit by leveraging the unique properties of negative triangularity plasmas on the DIII-D tokamak. Our findings suggest that the density limit, at least in this configuration, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit due to radiative edge cooling, and a core density limit driven by avalanche-like turbulent transport. Moreover, this study demonstrated sustained, non-disruptive operation at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
At the plasma edge, the density limit manifests as an abrupt saturation event. This limit is triggered by the onset of enhanced turbulence driven by a radiative instability on the high-field side. While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
In contrast, the core density is not directly constrained by the edge cooling and continues to rise well beyond the Greenwald limit. Its saturation is instead governed by a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, $1/f$-type power spectra, and elevated Hurst exponents. This highly persistent, non-diffusive transport regime ultimately prevents further increases in the core density.
The dual mechanisms of regional density saturation are consistent with the disparate power scalings observed for the core ($n_{\text{e,core}}\propto P_{\mathrm{SOL}}^{0.27}$) and edge ($n_{\text{e,edge}}\propto P_{\mathrm{SOL}}^{0.42}$). The power-dependent nature of this limit and its link to transport physics show a compelling resemblance to density limits observed in stellarators. This suggests a potentially universal physics of power balance and transport that transcends specific magnetic configurations, pointing to a more fundamental basis for the density limit in magnetically confined plasmas.
Furthermore, these results suggest that high Greenwald fraction operation is feasible in future tokamaks in negative triangularity configurations, provided that the edge cooling and core turbulence dynamics are well understood and controlled. Specifically, innovative divertor solutions should be developed to manage the edge radiative cooling, while advanced core turbulence control should be explored to optimize core density and confinement, thereby extending the operational space of future reactors.
Acknowledgments
The authors gratefully acknowledge the invaluable support and contributions of the entire DIII-D team in performing this experiment. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards Nos. DE-FC02-04ER54698, DE-SC0019352, DE-SC0019004, DE-FG02-97ER54415, DE-AC52-07NA27344, and DE-SC0016154. The author (P.H.D.) acknowledges support from LLNL-led SciDAC ABOUND Project SCW1832. The author (O.S.) performed this work within the framework of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No. 101052200– EUROfusion) and funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.
Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.

### v3

@@
-Disentangling Core and Edge Mechanisms of the L-mode Density Limit in DIII-D
+Disentangling Core and Edge Mechanisms of the Density Limit in DIII-D Negative Triangularity Plasmas
 Abstract
-The L-mode density limit is re-examined on the DIII-D tokamak using a negative triangularity configuration to study the underlying physics.
+The density limit is investigated in the DIII-D negative triangularity (NT) plasmas which lack a standard H-mode edge.
 We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms.
@@
 The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below $n_{\mathrm{G}}$.
-In contrast, the core density continues to rise until it saturates at a level featured by substantially enhanced core turbulence.
+In contrast, the core density continues to rise until it saturates, a state characterized by substantially enhanced core turbulence.
 Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis $\approx 6$), elevated Hurst exponents, and a $1/f$-type power spectrum.
-These findings demonstrate that the L-mode density limit is determined by a combination of edge radiative instabilities and core turbulent transport.
+These findings suggest that the density limit in the low-confinement regime is determined by a combination of edge radiative instabilities and core turbulent transport.
 This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
@@
 However, tokamak operation is constrained by an operational density limit [2, 3, 4].
-Exceeding this critical density leads to a rapid loss of confinement, often culminating a plasma termination.
+Exceeding this critical density leads to a rapid loss of confinement, often culminating in a plasma termination.
 Overcoming this long-standing obstacle is therefore a critical step toward realizing economically viable fusion energy.
@@
 This scaling is notably independent of heating power.
-This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power, particularly in the low-confinement (L-mode) regime.
-Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that even in L-mode, the achievable density limit may scale positively with increased heating power, potentially exceeding the conventional Greenwald density.
-Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physics based on power balance that transcends specific magnetic configurations.
+This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power.
+Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that the achievable density limit in the low-confinement (L-mode) regime may increase with heating power, potentially exceeding the conventional Greenwald density.
+Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physical basis rooted in power balance and transport that transcends specific magnetic configurations.
 Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
-The underlying physics of the L-mode density limit involves a complex interplay of several processes.
+The physics underpinning the L-mode density limit involves a complex interplay of several processes.
 Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors.
@@
 To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters.
-This avoids the complexities introduced by the L-H transition, edge transport barriers (ETBs), and edge localized modes (ELMs), which are characteristic of conventional H-mode plasmas.
+This avoids the complexities introduced by the edge transport barriers (ETBs) and edge localized modes (ELMs), which are characteristic of conventional high-confinement (H-mode) plasmas.
 The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig.
@@
 These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions.
-While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of a strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
-This paper presents a re-examination of the underlying physics governing the L-mode density limit using NT plasmas on the DIII-D tokamak.
+While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
+This paper presents an examination of the physics governing the density limit in NT plasmas on the DIII-D tokamak.
 We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power.
 A systematic power scan reveals disparate scaling dependencies for the core ($n_{e}\propto P^{0.27\pm 0.03}_{\text{SOL}}$) and the separatrix ($n_{e}\propto P^{0.42\pm 0.04}_{\text{SOL}}$) regions.
-Our detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas.
-At the plasma edge, the density is constrained substantially by enhanced turbulence driven by radiative instabilities and edge cooling.
+Detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas.
+At the plasma edge, the density is constrained by enhanced turbulence driven by a radiative instability.
 In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence.
@@
 II Experimental Arrangement
-The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power varied provided by neutral beam injection (NBI) and varied between discharges.
+The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power provided by neutral beam injection (NBI) and varied between discharges.
 All discharges employed a lower single-null configuration with an open divertor geometry, and the ion $\mathbf{B}\times\nabla\mathbf{B}$ drift was directed away from the primary divertor, as illustrated in Fig.
@@
 This paper focuses specifically on discharges with toroidal field $B_{t}=2.0$ T and plasma current $I_{p}=0.6$ MA.
-Additional experiments exploring different plasma current and toroidal field combinations will be reported separately [26].
+Additional experiments exploring different plasma current and toroidal field combinations are reported separately [26].
 In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber.
@@
 Figure 1: Cross-section of DIII-D tokamak showing locations of key diagnostics used for the present study.
-A comprehensive suite of diagnostics systems monitored the evolving plasma state, including its equilibrium profiles, radiated power, and turbulent fluctuations.
+A comprehensive suite of diagnostics monitored the evolving plasma state.
 Figure 1 shows the locations of the key diagnostics.
@@
 Several diagnostics were used to probe density fluctuations.
-A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along a horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side).
+A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along one horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side).
 The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane.
@@
 III.1.1 Density Saturation Beyond the Greenwald Limit
-An important result of this study is demonstration of the sustained L-mode plasma operation at densities exceeding the conventional Greenwald density limit.
+A key result of this study is the achievement of sustained L-mode operation at densities that surpass the conventional Greenwald density limit.
 This is shown in Fig.
@@
 The core density at $\rho\approx 0.7$, which is similar to the line-averaged density, consistently exceeds the Greenwald density ($n_{\text{G}}$).
-Crucially, the discharges with medium to high power NBI ($P_{\text{NBI}}>4$ MW) did not disrupt upon reaching their maximum density.
-Instead, they entered a phase of sustained saturated density that persisted until the termination of auxiliary heating.
+Notably, rather than disrupting upon reaching their maximum density, the discharges with medium to high NBI power ($P_{\mathrm{NBI}}>4~{}\mathrm{MW}$) transitioned into a phase of density saturation that persisted until the termination of auxiliary heating.
+The radiated power in the divertor [Fig.
+2(d)] and the main chamber [Fig.
+2(e)] also increase with density, indicating enhanced radiative losses.
+The spikes in the divertor radiation power [Fig.
+2(d)] mark the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
+The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge.
+For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26].
+This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density present during high-density operation.
+Therefore, Thomson scattering measurements are used for the analysis in this study.
 Figure 2: Time history of relevant plasma parameters in two representative NT discharges with different power input levels.
@@
 The dashed horizontal line in (b) and (c) indicates the Greenwald density.
-The radiated power in the divertor (Fig.
-2(d)) and the main chamber (Fig.
-2(e)) also increase with density initially, indicating enhanced radiative losses.
-The spikes in the divertor radiation power (Fig.
-2(d)) are associated with the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
-The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge.
-For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26].
-This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density that is presented during high-density operation.
-Therefore, Thomson scattering measurements are used for the analysis in this study.
-III.1.2 Distinct Power Scaling for Core and Edge Density Limit
-To explore the relationship between input power and achievable density, we examine the dependence of the maximum density on the power flowing into the scrape-off layer (SOL), $P_{\mathrm{SOL}}$.
-Here, $P_{\mathrm{SOL}}=P_{\text{in}}-P_{\text{rad,core}}-\frac{\mathrm{d}W}{\mathrm{dt}}$, accounting for total input power, core radiated power, and changes in stored energy.
-Averaged values over a 100 ms window are used.
+III.1.2 Disparate Power Scaling for Core and Edge Density Limit
+Having established that the density limit in these NT plasmas is a saturation phenomenon, we investigate its dependence on heating power.
+The analysis correlates the maximum achieved density with the power flowing into the scrape-off layer ($P_{\mathrm{SOL}}$), defined as $P_{\mathrm{SOL}}=P_{\mathrm{in}}-P_{\mathrm{rad,core}}-\mathrm{d}W/\mathrm{d}t$, accounting for total input power, core radiated power, and changes in stored energy.
+This parameter represents the effective supplied heating.
+All values are averaged over a 100 ms window at the point of maximum density.
+A systematic power scan reveals that the core and edge densities respond differently to $P_{\mathrm{SOL}}$, as shown in Fig.
+3.
+The maximum core density (at $\rho=0.7$) exhibits a weak but clear power dependence, scaling as $n_{e,\text{core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$.
+This result is consistent with L-mode density limit scalings reported on other tokamaks, such as ASDEX Upgrade [8].
+In contrast, the separatrix density displays a stronger relationship, scaling as $n_{e,\text{sep}}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$.
+This relatively strong power dependence is reminiscent of the density limit observed in stellarators, where power balance is known to be the dominant constraint [9, 10].
+These disparate power scalings constitute a central finding of this work; they indicate that the density saturation is not governed by a single, global mechanism.
+Instead, it suggests that distinct physical processes, with different sensitivities to heating power, are responsible for limiting the density in the core and edge regions.
+This hypothesis—that core saturation is dominated by turbulent transport while edge saturation is governed by radiative instabilities—guides the investigative framework for the remainder of this paper.
 Figure 3: The maximum density achieved at two different radial positions (core and separatrix) in negative triangularity (NT) discharges as a function of the loss power entering the scrape-off layer ($P_{\mathrm{SOL}}$).
 The horizontal dashed line indicates the Greenwald density.
-Figure 3 shows different scaling relationships for the core ($\rho=0.7$) and at the separatrix.
-The density at $\rho=0.7$ is chosen to represent the core region, as it is close to the line-averaged density measured by the Thomson scattering.
-The maximum core density increases with a relatively weak power dependence, scaling as $n_{e,\text{core}}\propto P^{0.27\pm 0.03}_{\mathrm{SOL}}$.
-This is consistent with L-mode density limit scalings observed on other devices such as ASDEX Upgrade [8].
-In contrast, the density near the separatrix exhibits a stronger dependence, scaling as $n_{e,\text{sep}}\propto P^{0.42\pm 0.04}_{\mathrm{SOL}}$.
-This power dependence is reminiscent of the power-dependent density limit observed in stellarators [9, 10], suggesting a more universal physical basis involving power balance and transport.
-These disparate scalings are a key finding, as they imply that distinct physical processes govern density saturation in the core and edge.
-The remainder of this paper will investigate these processes, proposing that turbulence dominates the core density limit, while radiative dynamics are critical at the edge.
-III.2 Temporal Evolution of Profiles Approaching the Density Limit
-III.2.1 Profiles Evolution Towards Density Limit
+III.2 Profile Evolution and Confinement Degradation Towards Density Limit
+III.2.1 Reduced Temperature and Rotational Shear at High Density
+As the plasma approaches the density limit, its equilibrium profiles evolve in a manner that signals a clear degradation of thermal confinement.
+This evolution, shown for a representative high-power discharge in Fig.
+4, provides important insight into the underlying transport dynamics.
 Figure 4: Temporal evolution of equilibrium profiles in a high-power discharge (shot 194133).
@@
 The dashed vertical line indicates the location of the separatrix.
-To investigate the underlying physical mechanisms, the temporal evolution of key plasma profiles is examined as each discharge approaches the density limit.
-Figure 4 displays the evolution of electron density ($n_{e}$), electron and ion temperature ($T_{e}$, $T_{i}$), and toroidal rotation rate ($\omega_{\phi}$) for a high-power discharge.
-As the overall density is ramped up (indicated by the color change from yellow to purple), several critical changes emerge.
-While the density profile (Fig.
-4(a)) rises across the plasma radius, it remains relatively flat in the core, indicating persistent particle transport.
-Concurrently, both the electron and ion temperature profiles (Figs.
-4(b) and 4(c)) decrease substantially.
-A prominent feature is the flattening of the $T_{e}$ profile at the edge, a classic sign of intensified edge cooling.
-The toroidal rotation profile (Fig.
-4(d)), which is initially peaked in the core, collapses as the density increases, leading to a significant reduction in core velocity shear.
+The electron density profile [Fig.
+4(a)] rises across the radius but remains notably flat in the core, indicating that strong particle transport persists even at high densities.
+The degraded confinement is accompanied by a loss of thermal energy, as both the electron and ion temperature profiles collapse [Figs.
+4(b), 4(c)].
+A noticeable feature is the pronounced flattening of the edge $T_{e}$ profile, a classic signature of intensified radiative cooling at the plasma boundary.
+Another critical change occurs in the plasma rotation.
+The toroidal rotation profile [Fig.
+4(d)], initially peaked in the core, progressively collapses with increasing density.
+This loss of core rotation directly dismantles the dominant component of the radial electric field ($E_{r}$), a key factor for turbulence suppression.
+As shown in Fig.
+5, the core $E_{r}$ shear, initially strong, diminishes and ultimately vanishes as the density limit is approached.
 Figure 5: The temporal evolution of the edge radial electric field ($E_{r}$) at the outer mid-plane in a high-power discharge (shot 194133).
@@
 The vertical dashed lines indicate the approximate location of the separatrix.
-Figure 5 shows the evolution of the edge radial electric field profile ($E_{r}$) at the outer mid-plane in the high-power discharge, derived from the carbon ion force balance equation.
-The $E_{r}$ profile initially peaks in the core region ($\rho<0.8$), but collapses progressively as the density increases.
-The edge $E_{r}$ profile remains relatively flat throughout the density ramp-up.
-This weakening of $E_{r}$ is primarily attributed to the reduced core toroidal rotation discussed previously.
-These profile changes point towards a degradation in plasma confinement.
-The collapse of the core rotation and the associated $E_{r}\times B$ velocity shear likely weakens the turbulence suppression.
-This loss of flow shear facilitates the enhanced turbulent transport that ultimately limits the core density, a topic explored in subsequent sections.
+This cascade of events, from intensified edge cooling to the collapse of core rotation and $E_{r}$ shear, indicates that the plasma’s ability to suppress turbulence deteriorates with rising density.
+The loss of this stabilizing flow shear enables enhanced turbulent transport that governs the core density limit, as subsequent sections will demonstrate.
 III.2.2 Divergent Evolution of Turbulence in Different Regions
+The degradation of thermal confinement described above coincides with a dramatic divergence in turbulent activity across the plasma, as illustrated in Fig.
+6 for a high-power discharge.
+A critical transition occurs at approximately 3700 ms (vertical dashed line), marking the onset of strong edge cooling.
+Before this event, the turbulence evolution is relatively uniform; afterward, it diverges sharply between different spatial regions.
 Figure 6: Time evolution of electron density and temperature and turbulence characteristics during a high-density discharge (shot 194133): (a) Local $n_{e}$ measured at $\rho$ = 0.7 and 0.95; (b) Local $T_{e}$ at $\rho$ = 0.7 and 0.95; (c) Spectrogram and (d) RMS amplitude of density fluctuations measured by Doppler back-scattering at the low-field side mid-plane ($\rho\approx 0.95$); (e) Spectrogram and (f) RMS amplitude of line-integrated density fluctuations measured by CO2 interferometer at the high field side; (g) Spectrogram and (h) RMS amplitude of line-integrated density fluctuations measured by RIP along the mid-plane.
 The vertical dashed line indicates the onset of enhanced turbulence.
-As the plasma approaches density saturation, a clear divergence in turbulent behavior emerges between different plasma regions.
-Figure 6 illustrates this by tracking electron density and temperature alongside multichannel turbulence measurements during a high-power discharge.
-The vertical dashed line at approximately 3700 ms marks the onset of this divergence, which features enhanced turbulence at the HFS and the core, likely due to radiative cooling.
-This critical event triggers distinct turbulence responses.
-Initially, the local density increases across the radius (Fig.
-6(a)).
-Concurrently, the local intermediate-$k$ turbulence at the low-field side (LFS) mid-plane ($\rho\approx 0.95$), measured by the Doppler backscattering systems, shows a steady increase in the root-mean-square (RMS) amplitude (Figs.
-6(e) and 6(f)).
-However, following the onset of the radiative cooling, a dramatic shift occurs: the edge density abruptly saturates (blue in Fig.
-6(a)), the electron temperature drops sharply (Fig.
-6(b)), and the LFS edge turbulence amplitude also saturates (Fig.
-6(d)).
-In contrast, line-integrated, low-$k$ density fluctuations at the HFS (Figs.
-6(e) and 6(f)) and midplane (Figs.
-6(g) and 6(h)) show a different evolution.
-The fluctuation amplitudes from both the HFS CO2 interferometer and the midplane RIP diagnostic begin to grow substantially after the onset of edge cooling.
-This divergent behavior—saturation of localized LFS turbulence contrasted with a continued, strong rise in line-integrated HFS and midplane fluctuations—points to a significant poloidal asymmetry in the transport.
-The turbulence becomes intense in the HFS region where the radiative front develops, linking the enhanced fluctuations and transport to the dynamics of the edge cooling potentially driven by radiative instabilities.
-III.3 The Edge Density Limit: Role of Radiative Cooling and Instabilities
-III.3.1 Development of Radiation Front at the HFS edge
-The evolution of profiles and turbulence suggests the formation of a radiative front and the resulting edge cooling as a potential limiting mechanism for the edge density.
-To elucidate the role of this radiative instability, the analysis now focuses on the edge region to examine the dynamics of the 2D radiated power distribution.
+Initially, both core and edge densities rise together [Fig.
+6(a)], accompanied by a steady increase in the amplitude of local, intermediate-$k$ turbulence at the low-field side (LFS) midplane ($\rho\approx 0.95$), as measured by Doppler backscattering [Fig.
+6(d)].
+This period of evolution ends abruptly at the 3700 ms transition.
+Immediately following this transition, the LFS turbulence saturates.
+This saturation coincides with the clamping of the edge density and the sharp drop in local electron temperature [Fig.
+6(b)].
+In contrast, turbulence in other regions intensifies.
+Line-integrated, low-$k$ density fluctuations on the high-field side (HFS), measured by the CO2 interferometer, grows substantially [Figs.
+6(e), 6(f)].
+A similar growth is observed in the line-integrated midplane fluctuations measured by the RIP diagnostic [Figs.
+6(g), 6(h)].
+This divergent evolution—saturation of localized turbulence on the LFS versus continued, strong growth of HFS and core fluctuations—reveals a profound poloidal asymmetry in the transport dynamics.
+This asymmetry strongly suggests that the density limit is not a globally uniform process but is instead triggered by a localized instability on the high-field side—likely the radiative phenomenon that we investigate next.
+III.3 The Edge Density Limit: A Radiatively Driven Phenomenon
+III.3.1 Visualizing the HFS Radiation Front
+The poloidally asymmetric turbulence described previously points to a localized radiative instability as the trigger for the edge density limit.
+Here, we examine the 2D radiated power distribution, reconstructed from bolometer measurements, to visualize the formation and evolution of this instability.
+Figure 7 presents a time-sequence of this evolution.
+While geometrically similar to a disruptive MARFE [16, 36], the feature observed here is a quasi-stable, non-disruptive radiation front that is directly linked to the observed density saturation.
 Figure 7: Temporal evolution of the 2D distribution of radiated power during a representative high-power discharge (shot 194133), reconstructed from bolometer measurements.
 (a) Before density ramp-up ($t=2800$ ms), (b) enhanced divertor radiation ($t=3800$ ms), (c) radiation front leaving inner divertor ($t=4800$ ms), and (d) radiation front intensified at the HFS ($t=5800$ ms).
-Figure 7 shows the temporal evolution of the radiated power distribution, reconstructed from bolometer measurements in a high-power discharge.
-This radiative feature is geometrically similar to a MARFE (Multifaceted Asymmetric Radiation From the Edge) [16, 36].
-However, in these experiments, it manifests as a non-disruptive, quasi-stable radiation front that leads to enhanced fluctuations and density saturation.
-The intense but localized radiative edge cooling has also been observed in divertor detachment experiments in NT discharges [37].
-Initially (Fig.
-7(a), $t=2800$ ms), before the density ramp-up, radiation is concentrated near the inner divertor target, typical of an attached divertor regime.
-As the density is raised (Fig.
-7(b), $t=3800$ ms), enhanced radiation emerges near the X-point, signaling the development of the radiative front.
-With a further density increase (Fig.
-7(c), $t=4800$ ms), the condensation front becomes fully developed and moves upward along the high-field side separatrix, a process strongly correlated with the edge cooling shown previously in Fig.
+The evolution of this front proceeds in clear stages.
+Early in the discharge ($t=2800$ ms), before the density ramp, the radiation is concentrated at the inner divertor target, characteristic of a standard attached divertor [Fig.
+7(a)].
+As plasma density increases ($t=3800$ ms), an intense radiation zone emerges near the X-point, signaling the birth of the radiative front [Fig.
+7(b)].
+This event coincides with the onset of the divergent turbulence seen in Fig.
 6.
-Finally, the radiation front propagates further towards the HFS (inboard), and the radiated power near the separatrix intensifies (Fig.
-7(d), $t=5800$ ms).
-It is noteworthy that in this discharge the disruption occurs shortly after a drop in the NBI power (at $t\approx 5800$ ms), which triggered an $n=1$ locked mode, and is not directly caused by the radiative edge cooling.
-III.3.2 Edge Density Saturation by Radiative Instabilities
-Previous observations suggest that the edge density saturates as radiative cooling intensifies.
-To quantify this relationship, we compare the evolution of the edge density with the amplitude of line-integrated density fluctuations.
-As the interferometer chord on the high-field side (HFS) passes directly through the developing radiation front, its signal serves as a reliable proxy for the intensity of the radiative instability.
+With a further increase in density ($t=4800$ ms), this front detaches from the divertor and propagates upward along the HFS separatrix [Fig.
+7(c)].
+This movement is strongly correlated with the localized edge cooling and turbulence saturation.
+Finally ($t=5800$ ms), the front moves further inboard and its emissivity intensifies [Fig.
+7(d)].
+This sequence provides direct visual evidence of a developing radiative instability on the HFS.
+This dynamic structure appears to be the primary driver of the edge density limit.
+Notably, the eventual discharge termination was caused by a subsequent drop in NBI power that triggered an $n=1$ locked mode and was not a direct consequence of the radiation front.
+III.3.2 Radiative Cooling and the Onset of HFS Turbulence
+The formation of the HFS radiation front actively cools the plasma and triggers a dramatic shift in the local turbulent state.
+Figure 8 provides quantitative evidence of this transition.
+As the HFS radiated power increases beyond a critical threshold, the edge plasma parameters and HFS turbulence changes fundamentally.
+The RMS amplitude of the line-integrated HFS density fluctuations ($\tilde{n}_{\mathrm{HFS}}$), measured by the CO2 interferometer passing through the radiation front, serves as a proxy for the instability’s strength.
+Figure 8: Transition to a degraded confinement state driven by HFS radiated power.
+(a) The edge adiabaticity parameter at $\rho\sim 0.9$ drops below unity as radiated power exceeds a critical threshold.
+(b) This transition coincides with a surge in the RMS amplitude of HFS density fluctuations ($\tilde{n}_{e}$).
+Data points are colored by the $H_{98y2}$ energy confinement factor, showing a clear shift from a high-confinement (yellow) to a low-confinement (purple) state.
+The plot is in log-log scale.
+20 ms moving average is applied to fluctuation data.
+The effect of edge cooling can be characterized by the dimensionless adiabaticity parameter [37], $\alpha_{\mathrm{adia}}=k^{2}_{\parallel}v^{2}_{te}/\omega\nu_{ei}\propto\frac{T_{e}^{5/2}}{n_{e}}$, which represents the comparison between parallel electron motion and turbulent timescales.
+Here, $k_{\parallel}$ is the parallel wavenumber of turbulence, $v_{te}$ is the electron thermal velocity, $\omega$ is the turbulence frequency, and $\nu_{ei}$ is the electron-ion collision frequency.
+When $\alpha_{\mathrm{adia}}<1$, the plasma enters a non-adiabatic regime where drift-wave instabilities are expected to grow.
+As shown in Fig.
+8(a), the onset of the strong HFS radiation drives $\alpha_{\mathrm{adia}}$ to plummet below 1.
+This transition into a non-adiabatic state coincides with a surge in HFS turbulence.
+Figure 8(b) shows that the RMS amplitude of HFS density fluctuations ($\tilde{n}_{e}$) increases sharply as the radiated power crosses the same threshold.
+The entire event is correlated with a degradation in plasma performance, as indicated by the drop in the $H_{98y2}$ confinement factor (color scale).
+This finding demonstrates that the HFS radiation front alters the local plasma state, unleashing the turbulence that may ultimately limit the edge density.
+III.3.3 Clamping of the Edge Density
+The preceding evidence indicates that the HFS radiative instability is the cause of the edge density saturation.
+To test this hypothesis quantitatively, we correlate the local edge density with the instability’s strength, again using the RMS amplitude of the line-integrated HFS density fluctuations ($\tilde{n}_{\mathrm{HFS}}$) as a proxy.
+Figure 9 plots the edge density (at $\rho=0.95$) against $\tilde{n}_{\mathrm{HFS}}$ on a log-log scale.
+The data, compiled from multiple discharges, reveals a dramatic bifurcation in the plasma behavior, defined by two distinct operational regimes that are distinguished by the energy confinement factor, $H_{98y2}$.
 Figure 9: Edge density at $\rho=0.95$ is plotted against the RMS amplitude of the density fluctuations measured by the HFS CO2 interferometer.
@@
 The dashed horizontal line represents the Greenwald density.
-Figure 8 plots the edge density (at $\rho=0.95$) against the RMS amplitude of density fluctuations on a log-log scale.
-The data are collected from multiple discharges and colored by the energy confinement factor $H_{98y2}$, reveal a sharp bifurcation between two distinct operational regimes.
-In the low-turbulence regime, characterized by high $H_{98y2}$ values, the edge density increases nearly linearly with HFS turbulence, following a power law of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{1.01}$ (gray line in Fig.
-8(a)).
-This indicates a phase where the increasing density turbulence does not yet limit the achievable edge density and plasma confinement.
-This relationship breaks down abruptly as the HFS turbulence exceeds a critical threshold, marking a transition to a high-turbulence regime.
-Here, the fluctuation amplitude, now dominated by the radiative instability itself, continues to grow by nearly an order of magnitude.
-In stark contrast, the edge density saturates completely, exhibiting a nearly flat scaling of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.1}$ (cyan line in Fig.
-8(b)).
-Critically, this density saturation is accompanied by a sharp degradation in confinement, with $H_{98y2}$ dropping significantly.
-The overall scaling relationship is $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.26}$.
-This analysis suggests that the edge density limit is not a gradual process, but a sharp transition triggered by enhanced turbulence, likely driven by a radiative instability and the associated edge cooling.
-This instability clamps the edge density at a level consistently below the Greenwald limit and can be related to the loss of the edge confinement.
-III.3.3 Turbulence Dynamics and the Breakdown of Zonal Flow Regulation
-While the preceding analysis points to the radiative instability as the primary limit on edge density, it is crucial to understand the interplay between the turbulence and the developing radiation front.
-For this purpose, we use 2D measurements from the BES diagnostic, which provides a detailed look at ion-scale density fluctuations.
-This analysis utilizes a lower-power discharge, where the radiative cooling occurs at lower density, allowing for higher-quality BES signals.
+At low turbulence amplitudes, the plasma is in a high-confinement state (high $H_{98y2}$, yellow points).
+In this phase, the edge density and turbulence grow together, following a nearly linear relationship ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{1.0}$).
+This represents a state where turbulence-driven transport has not yet begun to limit the density.
+This linear relationship breaks down abruptly once the turbulence level surpasses a critical threshold, at which point the system transitions to a degraded-confinement state (low $H_{98y2}$, purple points).
+While the instability amplitude continues to grow by nearly an order of magnitude, the edge density saturates almost completely, exhibiting a nearly flat scaling ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{0.1}$).
+This bifurcation provides strong evidence that the edge density limit is not a gradual process, but a sharp transition triggered by the radiative instability.
+Once initiated, the instability clamps the edge density at a value consistently below the Greenwald limit, leading to a significant increase in HFS turbulence.
+The apparent overall relationship across both regimes ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{0.23}$) also masks the sharpness of this critical transition.
+III.3.4 Breakdown of Zonal Flow Regulation
 Figure 10: Radial profiles of LFS edge turbulence characteristics during radiative cooling in a low-power NT discharge (shot 194131): (a) adiabaticity parameter, (b) normalized density fluctuation amplitude, (c) mean poloidal phase velocity, and (d) Reynolds stress $\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, (e) particle flux $\langle\tilde{n}\tilde{v}_{r}\rangle$, and (f) turbulent spreading flux $\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$.
@@
 Particularly, the profiles corresponding to the onset of the radiative cooling (about 3700 ms) are marked by dashed curves in purple.
-Figure 9 shows the evolution of key turbulence characteristics at the LFS edge as the radiative front forms (onset around 3700 ms, dashed purple curves).
-A fundamental change in the turbulence character is signaled by the dimensionless adiabaticity parameter [38, 13], $\alpha_{\text{adia}}=k_{\parallel}^{2}v_{te}^{2}/(\omega\nu_{ei})$.
-Here, $k_{\parallel}=1/qR$ is the parallel wavenumber, $v_{te}$ is the electron thermal velocity, $\nu_{ei}=\frac{\sqrt{2}e^{4}n_{i}\ln\Lambda}{12\pi^{3/2}\epsilon_{0}^{2}m_{e}^{1/2}T_{e}^{3/2}}$ is the electron-ion collision frequency, and $\omega$ is the turbulence frequency.
-As the edge cools, $\alpha_{\text{adia}}$ drops below unity, indicating a transition to a non-adiabatic regime, which is known to enhance turbulence and weaken zonal flow generation [13, 14].
-Consistent with this transition, the normalized density fluctuation amplitude ($\tilde{n}/\bar{n}$, Fig.
-9(b)) grows steadily.
-The mean poloidal phase velocity ($\bar{V}_{\theta}$, Fig.
-9(c)) remains relatively unchanged until the onset of radiative cooling.
-Crucially, the ability of the turbulence to self-regulate via zonal flows collapses.
-This is evident in the Reynolds stress ($\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig.
-9(d)), the primary drive for zonal flows.
-The Reynolds stress flattens, indicating that the nonlinear transfer of energy from micro-turbulence to stabilizing macroscopic flows, $\bar{V}_{\theta}\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, has been severely compromised.
-The consequence of this weakly regulated turbulence is an increase in outward transport.
-The turbulent particle flux ($\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig.
-9(e)) is directed outward and grows, while the turbulent spreading flux ($\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig.
-9(f)) indicates that turbulence intensity is spreading inward from the separatrix.
-These detailed measurements indicate a potential causal sequence: edge cooling drives a transition to a non-adiabatic turbulent state where the zonal flow self-regulation mechanism fails.
-This leads to enhanced outward transport, providing the micro-physical link between the onset of strong radiative cooling and the ultimate saturation of the edge density.
-III.4 The Core Density Limit: Role of Avalanche-like Turbulence
-While the edge radiative condensation directly limits the edge density, the core density continues to increase, even surpassing the Greenwald density after the radiative front has formed.
-This divergence strongly suggests that a different, core-localized mechanism is at play.
-We now present evidence that this mechanism is a fundamental change in the nature of core turbulent transport.
-III.4.1 Signatures of Avalanche-like Transport in the Core
-The collapse of stabilizing $E_{r}\times B$ shear (as shown in Fig.
-5) allows enhanced core turbulence.
-This is confirmed by examining line-integrated density fluctuations from the RIP diagnostic at the midplane.
-As the plasma approaches its density limit, the statistical character of this turbulence undergoes a significant change.
+The onset of the HFS radiative instability has global consequences, leading to a failure of turbulence self-regulation across the plasma edge.
+We use BES measurements on the low-field side to observe this breakdown in detail (Fig.
+10), revealing the microphysical mechanism responsible for enhanced transport.
+The trigger for this breakdown is the transition to the non-adiabatic regime.
+As the edge cools with the onset of the radiation front ($t\approx 3700$ ms), the dimensionless adiabaticity parameter, $\alpha_{\mathrm{adia}}$, drops below unity [Fig.
+10(a)].
+This transition is known to fundamentally alter drift-wave turbulence, weakening the mechanisms that generate stabilizing zonal flows [37, 13].
+While the fluctuation amplitude on the LFS saturates (as seen in Fig.
+6), the underlying dynamics show a flattening of the Reynolds stress profile [$\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig.
+10(d)].
+This collapse of the primary nonlinear drive for zonal flows (i.e., the Reynolds force, $\mathcal{F}_{\text{Re}}=-\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$) signifies a breakdown in the zonal flow regulation mechanism.
+Concurrently, the mean poloidal phase velocity of turbulence [$\langle\tilde{v}_{\theta}\rangle$, Fig.
+10(c)] also collapses.
+This breakdown in zonal flow regulation directly drives particle loss.
+The outward turbulent particle flux [$\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig.
+10(e)] grows significantly, while the turbulent spreading flux [$\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig.
+10(f)] indicates that turbulence intensity is propagating inward from the separatrix.
+These measurements suggest a compelling narrative: edge cooling initiates a transition to the non-adiabatic turbulent regime, which disables zonal flow regulation and unleashes enhanced transport.
+III.4 The Core Density Limit: A Transition to Avalanche-like Turbulence
+While the edge density saturates below the Greenwald limit, the core density continues to rise, often reaching much higher values than $n_{\text{G}}$.
+This decoupling implies that a separate, core-localized mechanism must govern the core density profile.
+Here, we present evidence that this mechanism is not a simple enhancement of diffusive transport, but a fundamental transition in its character to intermittent, avalanche-like behavior.
+This transition is enabled by the collapse of the stabilizing $E_{r}\times B$ shear, as shown previously in Fig.
+5.
+III.4.1 Statistical Signatures of Avalanche-like Transport,
+The statistical properties of the core density fluctuations, measured by the RIP diagnostic at the midplane, reveal three distinct signatures of this transition.
+First, the probability density function (PDF) of the fluctuation amplitudes evolves from Gaussian to heavy-tailed.
+Early in the discharge the PDF is Gaussian [yellow in Fig.
+11(a)], characteristic of diffusive transport.
+As the density limit is approached, however, the PDF develops pronounced, non-Gaussian tails [purple in Fig.
+11(a)].
+This signifies a shift to a transport regime dominated by large-amplitude, intermittent events.
+The growing intermittency is quantified by the kurtosis of the PDF, which rises from the Gaussian value of 3 to about 6 [Fig.
+11(b)].
 Figure 11: (a) Time evolution of the probability density function of the amplitudes of line-integrated density fluctuations measured by the RIP system at the midplane.
@@
 The horizontal dashed line presents the kurtosis value of a Gaussian distribution.
-Figure 10(a) shows the evolution of the probability density function (PDF) of these fluctuation amplitudes (high-pass filtered at $f>5$ kHz to avoid machine vibration noise).
-Early in the density ramp-up, the PDF is Gaussian (yellow), characteristic of diffusive, random-walk-like transport.
-However, as the density limit is approached—and particularly after the onset of the edge radiative cooling—the PDF develops pronounced “heavy tails” (purple).
-This indicates a transition to a state dominated by large-amplitude, intermittent transport events, a clear departure from simple diffusive behavior.
-This transition is quantified by the kurtosis of the PDF (Fig.
-10(b)), a measure of its “tailedness.” The kurtosis rises significantly after the onset of the radiative cooling and continues to increase, confirming the growing intermittency.
-Further evidence comes from the power spectrum of the fluctuations (Fig.
-11(a)).
-After the radiative cooling onset, a distinctive “$1/f$-type” power-law scaling emerges at low frequencies.
-This type of spectrum is a recognized signature of self-organized criticality (SOC) and avalanche-like dynamics [39, 40, 41], where small perturbations can trigger cascading transport events of all sizes.
+Second, the frequency spectrum of the turbulence develops a $1/f$-type power-law scaling.
+As the plasma evolves with rising density, a distinctive power-law scaling emerges at a wide range of frequencies [Fig.
+12(a)].
+Such a spectrum is a classic signature of self-organized criticality (SOC) and avalanche dynamics, where transport occurs via cascading events spanning a wide range of scales [38, 39, 40].
+Third, the transport acquires long-range temporal correlations, quantified by the Hurst exponent ($H$), a measure of the persistence in a time series [Fig.
+12(b)].
+The Hurst exponent rises from $H\approx 0.5$, the value for uncorrelated, random-walk events (diffusive transport), to $H\approx 0.8$ as the density limit is approached.
+A value of $H>0.5$ indicates persistent, correlated dynamics, another key feature of avalanche-like transport.
+This rise in $H$ is strongly correlated with a substantial decrease in the energy confinement factor $H_{98y2}$, linking the emergence of this non-diffusive transport regime to the degradation in plasma performance.
+Collectively, these three independent signatures—heavy-tailed PDFs, a $1/f$ power spectrum, and an elevated Hurst exponent—provide a consistent and compelling picture: core transport undergoes a fundamental transition from a diffusive to an intermittent, avalanche-like regime, which may be the mechanism for the core density limit.
 Figure 12: (a) Auto-power spectrum of the line-integrated midplane density fluctuations measured by RIP at different times.
@@
 The horizontal dashed line indicates a Hurst exponent of 0.5, corresponding to a random walk process.
-Finally, the transition to an avalanche-like state is supported by the Hurst exponent ($H$), a measure of long-range temporal correlations in the data (Fig.
-11(b)).
-The Hurst exponent rises from $H\approx 0.5$ (indicative of uncorrelated, random events) to $H\approx 0.8$ (indicative of persistent, correlated dynamics) as the density limit is approached.
-Consistently, the energy confinement factor $H_{98y2}$ also decreases substantially when the Hurst exponent $H$ begins to increase, indicating a loss of confinement associated with this transition.
-Collectively, the heavy-tailed PDFs, increased kurtosis, $1/f$ spectrum, and elevated Hurst exponent point to a consistent picture: core transport transitions from a diffusive to an intermittent, avalanche-like regime, which is ultimately responsible for limiting the core density.
 III.4.2 Core Density Saturation by Avalanche-like Turbulence
-The evidence for a transition to avalanche-like transport leads to a clear hypothesis: this enhanced turbulence is the primary factor limiting the core density.
-To test this hypothesis, we examine the direct relationship between the core density (at $\rho\approx 0.7$) and the amplitude of the midplane turbulence, as shown in Fig.
-12.
-The data from multiple discharges, plotted on a log-log scale and colored by the energy confinement factor ($H_{98y2}$), reveal a distinct bifurcation between two transport regimes.
-In the low-turbulence regime, corresponding to high confinement ($H_{98y2}>1.0$, yellow data points), the core density rises robustly with the turbulence amplitude, following a relatively strong power-law scaling of $n_{\text{e,core}}\propto\tilde{n}_{rms}^{0.49}$.
-This represents a phase where the core can still be effectively fueled, and the density increases despite the presence of turbulence.
+The emergence of avalanche-like dynamics provides a clear hypothesis for the core density limit: the large, intermittent transport events ultimately clamps the density profile.
+We test this hypothesis by correlating the core density (at $\rho\approx 0.7$) with the RMS amplitude of the midplane turbulence measured by RIP.
 Figure 13: Core density at $\rho=0.7$ is plotted against the RMS amplitude of the density fluctuations measured by the midplane RIP interferometers on a log-log scale (a) and linear scale (b).
@@
 Power law relationships for low- and high-turbulence regimes are indicated by the gray and cyan lines, respectively.
-However, as the avalanche-like transport intensifies and confinement degrades ($H_{98y2}<1.0$, purple data points), this relationship changes dramatically.
-The scaling becomes substantially weaker, flattening to $n_{\text{e,core}}\propto\tilde{n}_{\text{rms}}^{0.17}$.
-This transition to a weaker scaling indicates that the avalanche-like density fluctuations increase turbulent transport so effectively that they prevent significant increases in the core density.
-Notably, this saturation occurs at densities well above the Greenwald limit, reaching nearly twice the Greenwald density.
-The saturation of the core density is more clearly visible in the linear-scale plot in Fig.
-12(b).
-The overall response of the core density to the turbulence amplitude scales as $n_{\text{core}}\propto\tilde{n}_{\text{rms}}^{0.26}$, which also underscores the saturated behavior with increasing turbulence.
-The weak dependence on turbulence in this regime aligns well with the macroscopic power scaling of the core density limit ($n_{\mathrm{e,core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found earlier.
-This alignment, combined with the avalanche-like dynamics, suggests that the core density limit is governed by a fundamental transition in the core turbulent transport itself, which ultimately saturates the density profile.
+As shown in Fig. 13, data compiled from multiple discharges reveal a clear bifurcation between two transport regimes, distinguished by their confinement properties ($H_{98y2}$).
+In the low-turbulence, high-confinement regime ($H_{98y2}>1.0$, yellow points), which corresponds to the diffusive state, the core density rises robustly with increasing turbulence amplitude, following a strong power-law scaling of $n_{e,\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.49}$.
+This phase represents a state where fueling remains effective despite the presence of turbulence.
+This behavior changes dramatically as the transport transitions to the high-turbulence, low-confinement regime ($H_{98y2}<1.0$, purple points), which corresponds to the avalanche-like state.
+Here, the scaling weakens significantly, flattening to $n_{e,\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.17}$.
+This transition to a much weaker scaling demonstrates that the core fueling has become highly inefficient.
+The avalanche-like fluctuations effectively prevent any significant increases in density, leading to a saturation of the core density profile.
+This saturation occurs at densities far exceeding the Greenwald limit, approaching $2n_{\mathrm{G}}$.
+The weak overall scaling of density with turbulence ($n_{\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.26}$) is consistent with the weak macroscopic power scaling of the core density limit ($n_{e,\mathrm{core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found previously.
+This alignment confirms that the core density limit is governed not by simple power balance, but by a fundamental transition in the turbulent transport itself, from diffusive to avalanche-like, which saturates the density profile.
 IV Conclusion and Discussion
-This study has re-examined the L-mode density limit by leveraging the unique properties of negative triangularity plasmas on the DIII-D tokamak.
-Our findings suggest that the density limit, at least in this configuration, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit due to radiative edge cooling, and a core density limit driven by avalanche-like turbulent transport.
-Moreover, this study demonstrated sustained, non-disruptive operation at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
-At the plasma edge, the density limit manifests as an abrupt saturation event.
-This limit is triggered by the onset of enhanced turbulence driven by a radiative instability on the high-field side.
+This study re-examined the underlying physics of the density limit using negative triangularity plasmas on the DIII-D tokamak.
+Our findings reveal that the density limit, at least in NT plasmas without an H-mode edge, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit triggered by a radiative instability, and a core density limit driven by avalanche-like turbulent transport.
+Moreover, this study demonstrate sustained, non-disruptive operations at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
+At the plasma edge, the density limit manifests as an abrupt saturation event triggered by a radiative instability on the high-field side.
 While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
@@
 This is supported by signatures including heavy-tailed PDFs, $1/f$-type power spectra, and elevated Hurst exponents.
-This highly persistent, non-diffusive transport regime ultimately prevents further increases in the core density.
+This persistent, non-diffusive transport regime prevents further increases in the core density.
 The dual mechanisms of regional density saturation are consistent with the disparate power scalings observed for the core ($n_{\text{e,core}}\propto P_{\mathrm{SOL}}^{0.27}$) and edge ($n_{\text{e,edge}}\propto P_{\mathrm{SOL}}^{0.42}$).
@@
 This suggests a potentially universal physics of power balance and transport that transcends specific magnetic configurations, pointing to a more fundamental basis for the density limit in magnetically confined plasmas.
-Furthermore, these results suggest that high Greenwald fraction operation is feasible in future tokamaks in negative triangularity configurations, provided that the edge cooling and core turbulence dynamics are well understood and controlled.
-Specifically, innovative divertor solutions should be developed to manage the edge radiative cooling, while advanced core turbulence control should be explored to optimize core density and confinement, thereby extending the operational space of future reactors.
+These findings also present a potential strategy for achieving high Greenwald fraction operation in future tokamaks, particularly those utilizing negative triangularity configurations.
+Success hinges on addressing the distinct physical mechanisms governing the density limit in the core and edge.
+The edge radiative instability can be mitigated through advanced divertor designs and impurity control aimed at managing boundary heat fluxes.
+Concurrently, advanced techniques to suppress core turbulence, such as strong rotational shear flows and Shafranov shift, can be developed to optimize density profiles and confinement.
+By pursuing these parallel control strategies, the operational space of future reactors can be extended into stable, high-density regimes well beyond conventional scaling.
 Acknowledgments